OpenStudy (anonymous):

If a = 11 and c = 23 determine the vaue of A I will provide you with a right triangle.

OpenStudy (anonymous):



OpenStudy (anonymous):

do you mean the angle A

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sin A= a/c=11/23 =>A = sin^(-1) 11/23

OpenStudy (anonymous):

oh okay ty